### v2

# Understanding the Reasoning Ability of Language Models From the Perspective of Reasoning Paths Aggregation

Abstract
Pre-trained language models (LMs) are able to perform complex reasoning without explicit fine-tuning. To understand how pre-training with a next-token prediction objective contributes to the emergence of such reasoning capability, we propose that we can view an LM as deriving new conclusions by aggregating indirect reasoning paths seen at pre-training time. We found this perspective effective in two important cases of reasoning: logic reasoning with knowledge graphs (KGs) and math reasoning with math word problems (MWPs). More specifically, we formalize the reasoning paths as random walk paths on the knowledge/reasoning graphs. Analyses of learned LM distributions suggest that a weighted sum of relevant random walk path probabilities is a reasonable way to explain how LMs reason. Experiments and analysis on multiple KG and MWP datasets reveal the effect of training on random walk paths and suggest that augmenting unlabeled random walk reasoning paths can improve real-world multi-step reasoning performance.  [^1] We open source the code at https://github.com/WANGXinyiLinda/LM_random_walk.
Large language models
1 Introduction
Figure 1: We hypothesize that the pre-training corpus can be viewed as generated from random walks on a reasoning graph over world knowledge/concepts. With each node $s_{i}$ representing concepts, $p_{j}$ can be viewed as arguments that connect them. Then we hypothesize that a language model (LM) training on such a corpus can be viewed as reasoning by a weighted aggregation of random walk paths that connect the entities in interest. $P_{\text{LM}}$ denote the LM distribution while $P_{D}$ denotes the random walk probability from the pre-training corpus. $w_{i}^{1}$ denotes the weight assigned to the first random walk path by the LM for argument $p_{i}$, and $w_{i}^{2}$ denotes the weight assigned to the second random walk path.
Recently, pre-trained large language models (LLMs) (Touvron et al., 2023a, b; Brown et al., 2020) have demonstrated remarkable capabilities in performing intricate reasoning tasks (Kojima et al., 2022). These tasks include problem-solving with world knowledge (Hendrycks et al., 2020; Suzgun et al., 2022), logical reasoning (Pan et al., 2023), and solving mathematical problems (Cobbe et al., 2021; Hendrycks et al., 2021). These models are typically not explicitly fine-tuned to solve these tasks. Recent research (Jain et al., 2023) also suggests that the supervised fine-tuning process following pre-training only learns a wrapper on top of the already existing model capabilities, instead of learning new ones. It is intriguing to understand how next-token prediction pre-training contributes to the emergence of such reasoning capability. A better understanding of this matter can also inspire new pre-training/fine-tuning techniques to improve these important abilities of LLMs.
It is well-known that LLMs acquire emergent abilities through extensive pre-training (Wei et al., 2022a). In this paper, we focus on elucidating the emergence of reasoning ability — the capacity to draw novel conclusions from existing knowledge, which has been less studied. Many recent works also attempt to understand this phenomenon. Some works focus on understanding Transformers’ reasoning capability by construction (Liu et al., 2023; Chi et al., 2023; Feng et al., 2023). Others try to provide post hoc mechanistic explanations (Geiger et al., 2021; Wu et al., 2023; Hanna et al., 2023) or understanding inference time in-context learning reasoning (Li et al., 2023; Razeghi et al., 2022; Wang et al., 2023). Our study is more relevant to the line of work analyzing the contribution of pre-training data to LM reasoning (Bi et al., 2023; Chen et al., 2023; Xiao & Liu, 2023; Zhou et al., 2023; Ramesh et al., 2023).
In contrast to these works, we adopt a Bayesian view and try to understand why next-token-prediction pre-training can unlock LMs’ reasoning ability. More specifically, we hypothesize that LMs can aggregate the indirect reasoning paths seen at pre-training time, through the next-token-prediction training objective. In a real-world scenario, the reasoning path can be a piece of text argument connecting two concepts. We hypothesize that, at inference time, this enables an LM to jump from one concept to another during its reasoning process, which could be verbalized by generating chain-of-thought (CoT) solutions (Wei et al., 2022b), or silent without generating outputs.
Prystawski et al. (2023) propose a different hypothesis that localized structure on dependencies between variables in training data is important for LM reasoning, especially CoT reasoning. Our hypothesis implies a similar property of the pre-training data: when two concepts are related by a reasoning path, they are highly likely to cooccur in the data and thus form a graph-like localized structure. One drawback of Prystawski et al. (2023)’s work is that their experiments equate reasoning to conditional probability estimation of boolean variables with intermediate variables, which can be considered overly simplified compared to real-world reasoning processes. In our paper, we aim to produce a more realistic analysis of the effect of training data by closely examining two predominant types of reasoning: logical reasoning and mathematical reasoning. In these two reasoning scenarios, we first construct unsupervised random walk paths, which are used to (continually) pre-train the LM with next-token loss. Then we adopt the pre-trained LM to perform reasoning tasks on unseen examples.
For logical reasoning, we analyze a straightforward yet general reasoning scenario: reasoning over knowledge graphs. A knowledge graph (KG) stores facts in the form of triples $(e_{1},r,e_{2})$, where $e_{1}$ and $e_{2}$ represent entities connected by the relationship $r$. KGs can be incomplete, lacking certain relations between existing entities. These missing relations can typically be inferred from the known triples stored in the KG by employing logical rules. For instance, the relation (A, isGrandChildof, C) can be derived from the triples (A, isSonOf, B) and (B, isSonOf, C). We formalize a reasoning path as a random walk path on the KG, which enables us to accurately compute its probability. We show that an LM pre-trained from scratch on random walk paths generated from a given KG can accurately deduce missing relation connections. We also analyze the KL divergence between LM output distributions and weighted/unweighted sums of random walk path probabilities, which are variances of the classic path ranking algorithm (PRA) (Lao et al., 2011). Our analysis suggests that the LM distribution shares many similarities with aggregating the probabilities of possible random walk paths in a logical-rule-aware manner, and is usually superior to them.
For mathematical reasoning, we focus on a more complex case of reasoning: solving math word problems (MWPs). Since it is very challenging to pre-train an LM from scratch to perform well on MWPs, which require both math deduction and language understanding, we propose to continue training on a pre-trained base LM. Based on the insights obtained from the KG reasoning analysis, We propose to create random walk reasoning paths from existing CoT training data, and test the effectiveness of next-token-prediction training on these unlabeled reasoning paths. More specifically, we construct a reasoning graph by regarding the reasoning state at each CoT step as the graph node. Then we reorder and reconnect the existing CoT steps to form the random walk paths on the graph. Experiment results on three MWP datasets, GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021), show consistent improvement compared to vanilla supervised fine-tuning, and a similar effect of random walk path length as in the KG reasoning case is observed.
Our findings can be summarized as follows: (a) We show in both reasoning scenarios that our weighted random walk reasoning paths aggregation hypothesis is one (of many) valid ways to explain how LMs may gain their reasoning ability; (b) We show that LMs can utilize unlabeled reasoning paths highly efficiently and show the potential of incorporating the random walk idea to real-world (continue) pre-training.
2 Logical Reasoning
We first analyze a well-controlled case of logic reasoning, knowledge graph (KG) reasoning, by pre-training a small Transformer over random walk paths from KGs. The KL divergence between aggregated random walk path probabilities and LM distribution shows that LM is very close to a weighted aggregation. We also show that KL divergence reflects how LMs assign weights to logical rules. We find that there is usually an optimal random walk path length for training LMs. These observations support our reasoning paths aggregation hypothesis.
2.1 Problem setting
Consider a knowledge graph $\mathcal{G}=\{(e^{i}_{1},r^{i},e_{2}^{i})\}_{i=1}^{N}$ consisting of $N$ triples, such that the head entity $e^{i}_{1}$ and tail entity $e_{2}^{i}$ are related by $r^{i}$ for all $i$. Let $\mathcal{R}$ denote the set of all possible relations and $\mathcal{E}$ denote the set of all entities. Our goal is to predict a set of unseen triples $\mathcal{T}=\{(e^{j}_{1},r^{j},e_{2}^{j})\}_{j=1}^{m}$, $e^{j}_{1},e_{2}^{j}\in\mathcal{E}$, $r^{j}\in\mathcal{R}$, by training a Transformer based generative language model (LM) from scratch on the given knowledge graph $\mathcal{G}$. To translate a triple into a sentence (i.e. a sequence of tokens), We add each entity $e^{i}$ and relation $r^{i}$ as a new token (<e_i> and <r_i>) to the Transformer’s vocabulary and translate each triple into a three-token sentence “<e_i> <r_j> <e_k>.”. In this way, we avoid using any natural language thus no semantic meaning of the entity or relation name will affect the LM prediction.
2.2 Language Model Pre-training
We construct the training data by performing random walks on the given KG $\mathcal{G}$. More specifically, we randomly sample a start entity $e\sim U(\mathcal{E})$, where $U(\cdot)$ denotes the uniform distribution. Then we perform a random walk on $\mathcal{G}$ from $e$ by sampling the next node with $e^{\prime}\sim U(C(e))$, and stop at a maximum path length $L_{max}$.
Then we translate each triple into a sentence and concatenate all the sentences in the sampled random walk path to become a paragraph. The paragraphs are then concatenated together and separated by the special end-of-sequence token to form text chunks of the same length. The training loss function is the next-token prediction loss:
$\small\mathcal{L}_{LM}(\theta)=\sum_{\mathcal{D}}\sum_{t=1}^{T}\log\frac{\exp{(f_{\theta}(w_{t+1}|w_{1:t}))}}{\sum_{w\in\mathcal{V}}\exp{(f_{\theta}(w|w_{1:t}))}}$ (1)
Here, $\theta$ denotes the LM parameters  [^2] We use a randomly initialized GPT-2 model (Radford et al., 2019).. $w_{i}\in\mathcal{V}$ represents a token in the LM vocabulary $\mathcal{V}$, and $w_{1:T}$ is a token sequence in the training data $\mathcal{D}$, where $T$ is the length of a text chunk.
To test the reasoning ability of a pre-trained LM, we format the testing triples as sentence completion tasks. For example, the triple ($e_{1}$, $r$, $e_{2}$) will be translated to the prompt “<e_1> <r>) ”, and let the LM predict the next token, then verify the prediction with the ground truth $e_{2}$. Note that, here the raw LM output distribution is over all entities and relations. To make the LM distribution more well-defined and simplify the following analysis, we take the LM output logits over all entities and define the LM output distribution as:
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\frac{\exp{(f_{\theta}(e_{2}|e_{1},r))}}{\sum_{e\in\mathcal{E}}\exp{(f_{\theta}(e|e_{1},r))}}$ (2)
2.3 Random Walk Paths Aggregation
Recall that our hypothesis is LM can aggregate the reasoning paths seen at the pre-training time. In the KG setting, we can explicitly define how the reasoning/random walk paths are aggregated. Inspired by the classic path ranking algorithm PRA (Lao et al., 2011), we define the aggregation of random walk paths $P_{w}$ as the exponential of a weighted sum of the probabilities of all appropriate random walk paths connecting the two target entities. More specifically, we are interested in a distribution $P_{w}(e_{2}|e_{1},r)$ for unseen $(e_{1},r,e_{2})$ in the form of:
$\small P_{w}(e_{2}|e_{1},r)=\frac{\exp(S_{w}(e_{2}|e_{1},r)/T)}{\sum_{e\in\mathcal{E}}\exp(S_{w}(e|e_{1},r)/T)}$ (3)
Here $S_{w}(e_{2}|e_{1},r)$ is a score/logits of $e_{2}$. $T>0$ is a temperature to rescale the weighted logits $S_{w}$ so that it can match the scale of LM logits $f_{\theta}$  [^3] In practice, we take $T=0.01$., and that $P_{w}(e_{2}|e_{1},r)$ and $P_{\text{LM}}(e_{2}|e_{1},r)$ are more comparable. The score $S_{w}(e_{2}|e_{1},r)$ is defined to be a weighted sum of the probability of following all possible logical rules going from $e_{1}$ to $e_{2}$:
$\small S_{w}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}w_{r}(h)P(e_{2}|e_{1},h)$
Here $\mathcal{H}$ denotes the set of all possible logical rules, and $h\in\mathcal{H}$ is a specific logical rule. $w_{r}(h)$ is the weight assigned to rule $h$ when inferring relation $r$. For example, a rule for inferring the locatedIn relation can be $h:$ ($e_{1}$, neighborOf, $e_{3}$) $\wedge$ ($e_{3}$, locatedIn, $e_{2}$). Formally, for a target relation $r$, we consider logic rules with conjunctive form. $\forall\{e_{i}\}_{i=0}^{n}\subset\mathcal{E}$,
$\small(e_{0},r,e_{n})\leftarrow(e_{0},r_{1},e_{1})\wedge(e_{1},r_{2},e_{2})\wedge...\wedge(e_{n-1},r_{n},e_{n})$
where $(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. We abbreviate such rule by $h=[r_{1},r_{2},...,r_{n}]$. We can formalize the set of all possible logic rules by =$\mathcal{H}=\{[r_{1},r_{2},...,r_{n}]|n\geq 1,r_{i}\in\mathcal{R}\}$.
Then the probability of following a specific logic rule $h\in\mathcal{H}$ between $e_{1}$ and $e_{2}$ during the random walk would be the sum of the probability of all possible random walk paths from $e_{1}$ to $e_{2}$ following the rule $h=[r_{1},r_{2},...,r_{n}]$:
$\small P(e_{n}|e_{0},h)=\sum_{(e_{0},r_{1},e_{1})...(e_{n-1},r_{n},e_{n})\in\mathcal{P}_{h}}\prod_{i=1}^{n}P(e_{i}|e_{i-1},r_{i})$
where $\mathcal{P}_{h}$ denotes all paths from the KG following $h$. Following the pre-training data generation, we perform a uniform random walk. i.e. $P(e_{i}|e_{i-1},r_{i})=1/|C(e_{i-1})|$. Then the rule probability $P(e_{2}|e_{1},h)$ can be computed directly from the KG.
To learn the rule weights $w_{r}$, we first observe that
$\small P_{w}(e_{2}|e_{1},r)=\frac{P_{w}(r|e_{1},e_{2})}{\sum_{e\in\mathcal{E}}P_{w}(r|e_{1},e)},$
if we sample $e_{1}$ and $e_{2}$ independently and uniformly. Recall Equation 3, we can instead model $P_{w}(r|e_{1},e_{2})\propto\exp S_{w}(e_{2}|e_{1},r)$. We can even further simplify it into a binary classification problem $p_{i}=P_{w}(\mathbbm{1}_{r^{i}=r}|e_{1}^{i},e_{2}^{i})$. Then we can use $w_{r}$ to parameterize a logistic regression model with a loss function:
$\small\mathcal{L}_{r}(w)=-\sum_{i}\left[y_{i}\ln{p_{i}}+(1-y_{i})\ln{(1-p_{i})}\right]+\lambda|w|,$
where $p_{i}=\frac{\exp{S_{w}(e_{2}^{i}|e_{1}^{i},r)}}{1+\exp{S_{w}(e_{2}^{i}|e_{1}^{i},r)}}$, and the binary label $y_{i}=\mathbbm{1}_{r^{i}=r}$. $\lambda|w|$ is a regularization term, and we can take any appropriate norm on $w$.
At training time, we sample positive triples with relation $r$ and negative triples with other relations from $\mathcal{G}$ as training data. We search over the graph to compute their probability of being reached by each rule $P(e_{2}|e_{1},h)$ to compute $p_{i}$.
For computation efficiency, we only want to search for a subset of more possible logical reasoning rules $\mathcal{H}_{r}$ in the test set for each relation $r$, and assign $w_{r}(h)=0$ for $h\notin\mathcal{H}_{r}$. Note that a rule can be infinitely long, so we set a maximum rule length $n\leq N_{max}$. To obtain $\mathcal{H}_{r}$, we search over $\mathcal{G}$, and record all paths between any two entities that are connected with the relation $r$, and shorter than $N_{max}$. We then collect the rules that have more than $m$ valid paths.
A simplified version of $P_{w}$ would be letting $w_{r}(h)=1$ for all $h$ and $r$. And we define this unweighted aggregation distribution to be $P_{s}$:
$\small P_{s}(e_{2}|e_{1},r)=\frac{\exp(\sum_{h\in\mathcal{H}_{r}}P(e_{2}|e_{1},h)/T)}{\sum_{e\in\mathcal{E}}\exp(\sum_{h\in\mathcal{H}_{r}}P(e|e_{1},h)/T)}$ (4)
Figure 2: KL divergence between various reference distributions and LM distribution, with different maximum random walk lengths, averaged over Countries (top) and UMLS (bottom) testing set, respectively. The rows correspond to the LM distribution $P_{\text{LM}}(e_{2}|e_{1},r)$ with maximum pre-training random walk path lengths ($L_{max}$) ranging from 1 to 10. From left to right, the columns correspond to the weighted aggregation distribution $P_{w}(e_{2}|e_{1},r)$ with maximum random walk path lengths ($N_{max}$) from 1 to 10, the unweighted aggregation distribution $P_{s}(e_{2}|e_{1},r)$ with maximum random walk path lengths ($N_{max}$) from 1 to 10, the reference distribution $P^{*}(e_{2}|e_{1},r)$, and the uniform distribution $P_{u}(e_{2})$, respectively. A darker color represents a smaller KL value, meaning that the two distributions are closer. In general, $KL[P_{w},P_{\text{LM}}]$ is always smaller than $KL[P_{s},P_{\text{LM}}]$, which implies that LM is learning the difference in rule importance. $KL[P^{*},P_{\text{LM}}]$ and $KL[P_{u},P_{\text{LM}}]$ serve as anchor points to show the scale of KL values. $KL[P^{*},P_{\text{LM}}]$ is generally high because the probability mass concentrates on correct answers, thus it can be very different from the LM distribution. Thus $KL[P^{*},P_{\text{LM}}]$ shows how peaky the LM distribution is, and $KL[P_{u},P_{\text{LM}}]$ shows how flat the LM distribution is.
Figure 3: Testing accuracy w.r.t. various maximum pre-training random walk lengths ($1\leq L_{max}\leq 10$) on Countries (left) and UMLS (right) datasets, respectively. For Countries, the LM ($P_{\text{LM}}$) performance converges to the weighted aggregation ($P_{w}$) performance, while for UMLS, LM consistently outperforms both weighted ($P_{w}$) and unweighted ($P_{s}$) aggregation performance. This is likely because LM ($P_{\text{LM}}$) can learn a better logical rule weighting scheme than weighted aggregation ($P_{w}$) in more complex KGs.
2.4 KL Divergence and Prediction Accuracy
To better understand the similarity between LM and the random walk aggregation algorithm as described in the previous section, we propose to compute and analyze the KL divergence between them: $KL[P_{w}(\mathbf{e}|e_{1},r),P_{\text{LM}}(\mathbf{e}|e_{1},r)]$, where $\mathbf{e}$ is a random variable taking values in $\mathcal{E}$. To better understand the meaning of the computed KL divergence, we derive an upper bound of it by writing $P_{\text{LM}}(e_{2}|e_{1},r)$ as marginalization over rules:
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$ (5)
Similarly, we can write
$\small P_{w}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)$ (6)
Then by the Log sum inequality, we can see that the KL divergence of the rule importance is an upper bound of the computed KL divergence  [^4] Proof available in Appendix.:
Proposition 2.1.
If LM effectively learned the random walk data distribution through pre-training, we have
[…]
Here $\mathbf{h}$ is a random variable taking values in $\mathcal{H}$. This means the KL divergence reflects how LM assigns probabilities to possible logical rules based on the given prompt, which implies how the LM learns to do logical reasoning.
KL computation We compute the KL divergence between the weighted aggregation distribution $P_{w}(e_{2}|e_{1},r)$ as defined in Equation 3 and the LM distribution $P_{\text{LM}}(e_{2}|e_{1},r)$ as defined in Equation 2, abbreviated as $KL[P_{w},P_{\text{LM}}]$. We then compare it with the KL divergence between the unweighted aggregation distribution $P_{s}(e_{2}|e_{1},r)$ as defined in Equation 4 and the LM distribution, abbreviated as $KL[P_{s},P_{\text{LM}}]$. To better understand the effect of random walk length, we consider maximum random walk path length ranging from 1 to 10 (i.e. $1\leq L_{max}\leq 10$ and $1\leq N_{max}\leq 10$), for computing both the aggregation distribution and the LM distribution. We then compute a pairwise KL between each of them and show the results as a heatmap. To better anchor the computed KL divergence, we also compute the KL divergence $KL[P^{*},P_{\text{LM}}]$ between a reference distribution $P^{*}$ and the LM distribution $P_{\text{LM}}$, and KL divergence $KL[P_{u},P_{\text{LM}}]$ between the uniform distribution $P_{u}$ and the LM distribution $P_{\text{LM}}$. Here $P^{*}$ is uniform over all correct answers, and $P_{u}$ is uniform over all possible answers. The described KL divergences for Countries (top) and UMLS (bottom) testing sets are shown in heatmaps in Figure 2. More interpretations of these quantities can be found in the caption.
Accuracy We also compute the prediction accuracy using each method and plot it w.r.t to path length ($1\leq L_{max}\leq 10$). Note that there could be more than one correct answer for a query $(e_{1},r)$. We say the prediction is correct as long as it is one of the correct answers. The described testing accuracy for Countries (left) and UMLS (right) is shown in Figure 3, where LM is $\arg\max P_{\text{LM}}$, Weighted is $\arg\max P_{w}$, and Unweighted is $\arg\max P_{s}$. In general, LM predictor $P_{\text{LM}}$ performs on par/better than weighted aggregation $P_{w}$, and significantly better than the unweighted aggregation $P_{s}$. This shows that LM likely learns a better logical rule weighting scheme than $P_{w}$.
2.5 Results and Analysis
Figure 4: Testing accuracy of LM trained on different random walk path lengths. Each line corresponds to a different KG dataset and thus is not directly comparable. We want to highlight the common trend here that each line peaks at some optimal path length.
We consider five KG datasets in total: Countries (Bouchard et al., 2015), UMLS (Kok & Domingos, 2007), Kinship (Denham, 2020), NELL-995 (Xiong et al., 2017), and FB15K-237 (Toutanova et al., 2015)  [^5] Dataset details can be found in Appendix.. We take the smallest two for KL divergence analysis for their lower time complexity. We show LM prediction accuracy for all datasets with different pre-training path lengths.
KL divergence with Countries In Figure 2 (top), we can see that when the maximum path length for computing the aggregated distribution (columns) is three, there is a sudden drop in $KL[P_{w},P_{\text{LM}}]$. This is because the ground truth path length to reach the correct answers in the testing set is three (fixed when constructing the dataset). Both the weighted and unweighted aggregation of random walk paths have low accuracy with path lengths less than three as shown in Figure 3 (left). The behavior of the path aggregation method is not well-defined at this stage and thus can result in an abnormal KL trend. On the other hand, LM yields a non-trivial accuracy when trained with a path length smaller than three, which shows LM’s ability to generalize beyond the pre-training reasoning length. This echoes the findings in Xiao & Liu (2023); Zhou et al. (2023), that Transformers can generalize to longer sequences than training sequences.
As shown in Figure 2 (top left), the weighted aggregation scheme $P_{w}$ converges to a stable distribution, likely by putting most weights on shorter rules when using long random walk paths. The LM distribution $P_{\text{LM}}$ becomes closer to $P_{w}$ when the pre-training path length becomes longer. On the other hand, $KL[P_{s},P_{\text{LM}}]$ stably increases when the path length for $P_{s}$ becomes larger. This echoes the accuracy trends as shown in Figure 3 (left). For the countries dataset, since it only has two relations, longer random walk paths introduce more noise than useful information. Thus by increasing the path length the unweighted aggregation scheme $P_{s}$ becomes less and less effective. Both $P_{w}$ and $P_{\text{LM}}$ learn to assign a small weight to the long/noisy paths, and thus do not experience an accuracy drop.
KL divergence with UMLS In Figure 2 (bottom), we can see that when the maximum path length for computing the aggregated distribution (columns) is larger than 3, the weighted aggregation scheme $P_{w}$ also converges to a stable distribution. To investigate why path length 3 is unique, we find the average path length corresponding to the largest number of valid paths for each relation in the testing set is 3.14. We find the average path length corresponding to the largest weight assigned by $P_{w}$ when $N_{max}=10$ is 2.75. This confirms that path length three is likely a good rule length for many relations. However, from Figure 3 (right), we can see that both weighted ($P_{w}$) and unweighted ($P_{s}$) aggregation peaked at path length two instead of three. We believe this is because when the rule length becomes larger (i.e. larger than two), the validity of a rule would be more head entity ($e_{1}$) dependent. Using only relation-dependent weight $w_{r}(h)$ as in $P_{w}$ is likely insufficient. This also explains why LM constantly outperforms both path aggregation methods: LM likely learns a rule importance function that depends both on the head entity and the relation.
Different from the Countries dataset, UMLS’ $KL[P_{s},P_{\text{LM}}]$ does not increase when the path length for $P_{s}$ increases. Instead, $KL[P_{s},P_{\text{LM}}]$ follows a similar trend as $KL[P_{w},P_{\text{LM}}]$, while in general $KL[P_{w},P_{\text{LM}}]$ is smaller than $KL[P_{s},P_{\text{LM}}]$. Similarly, in Figure 3 (right), the weighted ($P_{w}$) and unweighted ($P_{s}$) aggregation has a similar performance, while $P_{w}$ is slightly better. This shows that the logical rule weights learned by $P_{w}$ are similar between different rules, so it has similar effects (KL and accuracy) as the unweighted version $P_{s}$. The LM also has a flatter distribution, as we can see for UMLS $KL[P^{*},P_{\text{LM}}]<KL[P_{u},P_{\text{LM}}]$ while for Countries $KL[P^{*},P_{\text{LM}}]>KL[P_{u},P_{\text{LM}}]$. This is likely because UMLS is more complex than Countries (49 v.s. 2 relations), thus many longer paths and rules are similarly useful for prediction, making the LM distribution flatter.
Prediction accuracy v.s. pre-training path length We briefly touched on how the pre-training random walk path length $L_{max}$ affects the LM distribution in the analysis above. In general, a longer path length improves the prediction accuracy and decreases $KL[P_{w},P_{\text{LM}}]$. This shows that LM can improve the logical rule weight assignment when trained with a longer path length. To further investigate this problem, we pre-train LM on longer random walk path lengths with more KG datasets.
In Figure 4, we show the LM prediction accuracy v.s. the maximum pre-training random path length of 1, 5, 7, 10, 15, and 20, trained on five different KG datasets. In general, there is a large performance gain from a path length of 1 to 5. Note that when the path length is equal to one, we randomly sample individual triples from a KG. i.e. There are no reasoning paths in the training data. So it is important to have reasoning paths with a non-trivial length in the pre-training data, to enable the LM’s reasoning ability. By extending the maximum length from 10 to 20, we can see that there is a slight drop in the Countries dataset. Similarly, in most datasets, there is a small decrease after an optimal path length. This is likely because a too-long random walk path would contain more noise/unrelated triples for reasoning. i.e. It is less likely to be useful for predicting the head and tail entity relation in a path aggregation sense. On the other hand, we can understand this from a localized data structure perspective (Prystawski et al., 2023): a sufficiently long random walk path makes any two entities similarly possible to appear in the same path, thus hurting the local dependency in the training data.
3 Math Reasoning
After carefully analyzing the logical reasoning on KGs, we want to apply and verify the obtained insights on a more interesting and realistic case of reasoning: math reasoning with textual descriptions, which is usually called math word problems (MWPs). We continue training a pre-trained LM with random walk reasoning paths and show that these unlabeled paths consistently benefit math reasoning performance on three datasets. We also observe a similar optimal random walk path length effect as in the logical reasoning case, which is associated with the intrinsic reasoning length of different datasets. These results support our reasoning path aggregation hypothesis and imply principles for constructing/augmenting pre-training data.
3.1 Problem Setting
Suppose we have a set of training data $\mathcal{D}=\{(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}},y^{i})\}_{i}$, where $x^{i}$ is a question described in the text that needs to be solved by math technique. $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ is a chain-of-thought (CoT) solution, where $r^{i}_{j}$ is one reasoning step. $y^{i}$ is the ground truth answer to the question. Since MWP datasets are hard to collect and usually small in size, a model is not likely to generalize to new questions by aggregating reasoning paths over this small set of CoT reasoning paths. Fine-tuning on a pre-trained LM can effectively mitigate this problem since the LM has already seen many other reasoning paths at the pre-training time, but more unlabeled reasoning paths specific to this task would likely improve the testing performance if the path aggregation hypothesis still holds for this task.
3.2 Random Walk on Latent Reasoning Graph
We assume that CoT paths $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ can be regarded as random walk paths sampled from a reasoning graph $\mathcal{G}$, where the nodes are the reasoning states at each step $r^{i}_{j}$. The reasoning state can be regarded as a belief that will be updated after each reasoning step. Denote the last hidden state pre-trained LM we are going to tune by $f_{\theta}$. To represent the reasoning state for each step $r^{i}_{j}$, we propose to use $f_{\theta}$ to cumulatively encode all the steps before $r^{i}_{j}$, and then average over the sequence dimension, to obtain a fixed dimensional vector $s^{i}_{j}$:
$\small s^{i}_{j}=\text{avg}\;f_{\theta}(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{j})$
Assuming similar $s^{i}_{j}$’s are sampled from the same node of the latent reasoning graph, we propose to cluster  [^6] In practice we use K-meanings clustering. similar $s^{i}_{j}$’s together to form a node. Suppose we have constructed a graph $\mathcal{G}$ from the CoT dataset $\mathcal{D}$, with nodes $A_{1},A_{2},...,A_{K}$, where $K$ is predefined by the clustering algorithm. Each CoT step would be classified into a node. i.e. $r^{i}_{j}\in A_{m}$ for some $m\in[1,k]$. Then we can perform random walks on the graph by using the original CoT as links between the nodes as shown in Algorithm 1. Then we record the random walk paths produced by Algorithm 1 and do next-token-prediction training on them for $M$ steps. To make sure the LM can produce a CoT solution and a final answer, we do another $N-M$ step of supervised fine-tuning (SFT) on the original dataset $\mathcal{D}$, for some $N>M$.
Algorithm 1 Random Walk on Latent Graph
  Input: CoT dataset $\mathcal{D}$, latent graph $\mathcal{G}$, maximum path length $L_{max}$.
  Randomly initialize current node $a=A_{k}$. Initialize path $p=[]$
  repeat
    Randomly choose a CoT step $r^{i}_{j}\in a$.
    Uniformly sample $m$ from $[1,L]$.
    Append $r^{i}_{j},r^{i}_{j+1},...,r^{i}_{\min\{j+m,n^{i}\}}$ to path $p$.
    Suppose $r^{i}_{\min\{j+m,n^{i}\}}\in A_{l}$. Set $a=A_{l}$.
  until $\text{len}(p)\geq L_{max}$.
3.3 Experiments
Datasets. We conduct experiments on three math word problem (MWP) datasets: GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021).  [^7] Dataset details can be found in Appendix.
Training We do LoRA (Hu et al., 2021) parameter efficient training in 8 bits with Llama 2 models (Touvron et al., 2023b) because of computation limits. If not specified, we default to using the 7B version.
Results. In Table 1, we demonstrate the effectiveness of our proposed method against the supervised fine-tuning (SFT) baseline. We train both our method and SFT with $N=2500$ steps in total. The first $M=500$ steps of our method are continually pre-trained on random walk data, and then we do 2000 steps of SFT on the original dataset. Our method is about 3% better in accuracy on average over three MWP datasets and two Llama 2 models.
Then we investigate the effect of random walk path length $L_{max}$ by plotting accuracy v.s. path lengths. In Figure 5, we observe that each dataset has a performance peak at a certain random walk length. While both AQUA and GSM8K peak at path length 10, the SVAMP dataset peaks at path length 5. This is likely related to the different intrinsic reasoning lengths for different datasets. The average length of CoTs in AQUA, GSM8, and SVAMP training sets are 4.79, 3.72, and 1.36, respectively. The reasoning length required for SVAMP is significantly shorter than the other two datasets, thus explaining the earlier peaking. As we discussed in the logical reasoning case, a long random walk may introduce more noise than useful information. Note that even the LM performance can drop after the optimal path length, it is always better than training with path length one. i.e. multi-step random walk always helps.
We also do ablation studies on two critical hyperparameters of our method: the number of steps training on random walk paths $M$ and the number of clusters/nodes $K$. In Table 2, we show that the optimal number of training steps $M$ is 500 for all three datasets. Since the generated random walk reasoning paths are not natural within small corpora, e.g. the subject might be suddenly changed from one step to another, training too many steps might make the LM overfit the unwanted artifacts. In Table 3, we show that the optimal number of clusters is 100 for all three datasets. Here 0 clusters mean the SFT baseline. Since the datasets we use are small in scale, clustering with a large number of clusters may introduce more noise than useful matchings. We hypothesize that this may be solved by using a larger dataset and more number of clusters/nodes $K$: in this case, the steps within each node will be more intrinsically similar. This also hints at the potential of our method in the actual pre-training stage: we can view each example in the pre-training corpus as a reasoning path and apply our method.
Model Method GSM8K AQUA SVAMP Avg.
7B SFT 26.8 30.0 53.3 36.7
Ours 28.5 34.6 55.8 39.6
13B SFT 37.1 35.0 66.4 46.2
Ours 41.2 37.4 69.0 49.2
Table 1: Testing accuracy of different size Llama 2 models continue pre-trained with our random walk paths and then supervised fine-tuned. The supervised fine-tuning baseline (SFT) is fine-tuned by the same number of total steps. Results are reported on three math word problem (MWP) datasets.
Figure 5: Testing accuracy of continue pre-training with our random walk paths of different length $L_{max}$. Each line corresponds to a different MWP dataset and thus is not directly comparable. We want to highlight the common trend here that each line would peak at some optimal path length range, which is similar to Figure 4.
#Steps GSM8K AQUA SVAMP Avg.
0 26.8 30.0 53.3 36.7
200 27.5 30.1 53.6 37.1
500 28.5 34.6 55.8 39.6
1000 24.9 32.3 51.6 36.3
Table 2: Ablation on the number of random walk training steps $M$.
#Nodes GSM8K AQUA SVAMP Avg.
0 26.8 30.0 53.3 36.7
10 26.8 30.3 54.8 37.3
50 26.6 29.9 54.7 37.1
100 28.5 34.6 55.8 39.6
200 26.6 31.1 52.5 36.7
Table 3: Ablation on the number of clusters/nodes $K$.
4 Related Work
Many recent works are investigating LM’s reasoning ability. Geiger et al. (2021); Wu et al. (2023) ai to find the causal abstraction of an LM. (Hanna et al., 2023) tries to find circuit for year-span-prediction. Liu et al. (2023); Chi et al. (2023); Feng et al. (2023) show that CoTs enable fixed-size Transformers to perform certain types of reasoning tasks. Li et al. (2023); Razeghi et al. (2022); Wang et al. (2023) try to understand inference time in-context CoT reasoning. Our study is more relevant to the line of work analyzing the contribution of pre-training data to LM reasoning. Bi et al. (2023) analyzes how code data affect program-of-thoughts (Chen et al., 2023) reasoning ability. Xiao & Liu (2023); Zhou et al. (2023) study how reasoning length generalizes from training data. Ramesh et al. (2023) studies LMs’ compositional generalization ability. Our hypothesis also echos the conclusion of Malach (2023) that reasoning paths in training data enable supervision on intermediate steps with a next-token-prediction objective. Prystawski et al. (2023) propose a different hypothesis that localized structure on dependencies between variables in training data is important for LM reasoning, especially CoT reasoning. Our proposed hypothesis echoes theirs and is shown effective on more realistic data and tasks. [^8] More related work on logical reasoning and math reasoning can be found in the Appendix.
5 Conclusion
In conclusion, we aim to understand reasoning abilities in language models (LMs), from the perspective of aggregating reasoning paths from pre-training data. The findings shed light on the origins of LLMs’ remarkable reasoning capabilities, showcasing the importance of pre-training in acquiring these skills. The construction of the pre-training sequence, such as organizing it as ”chains” or random walks on the graph, was found to significantly impact the effectiveness of reasoning. The study also revealed that LM behavior is similar to reason over known facts by aggregating relevant reasoning paths. These insights contribute to our understanding of the underlying mechanisms behind LLMs’ reasoning abilities and lead to a potential pre-training data augmentation technique to boost reasoning performance.
Impact Statement
Understanding the reasoning processes of large language models (LLMs) through the lens of aggregating indirect reasoning paths holds potential implications for identifying and mitigating potential biases within LLMs. By formalizing reasoning as random walk paths on knowledge and reasoning graphs, this approach not only elucidates the mechanisms through which LLMs derive conclusions but also sheds light on data and reasoning paths that contribute to their outputs. This insight is crucial for recognizing biases embedded in the training data or in the reasoning process itself. Recognizing these biases is the first step toward developing more equitable and transparent models. By augmenting models with unbiased, unlabeled random walk reasoning paths, we can potentially reduce the influence of biased reasoning patterns and improve the fairness and reliability of LLMs in real-world applications. This research advances our understanding of LLM reasoning capabilities and their implications for bias, paving the way for more responsible AI development and deployment.
Appendix A Proof
Proposition A.1.
If LM effectively learned the random walk data distribution through pre-training, we have
[…]
Proof.
Recall that
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$
and
$\small P_{w}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r).$
By log sum inequality, we have:
$\text{KL}[P_{w}(\mathbf{e}|e_{1},r),P_{\text{LM}}(\mathbf{e}|e_{1},r)]$
$=$ $\sum_{e_{2}\in\mathcal{E}}P_{w}(e_{2}|e_{1},r)\log\frac{P_{w}(e_{2}|e_{1},r)}{P_{\text{LM}}(e_{2}|e_{1},r)}$
$\leq$ $\sum_{e_{2}\in\mathcal{E}}\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)\frac{P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)}{P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)}$
$=$ $\sum_{e_{2}\in\mathcal{E}}\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)\frac{P_{\text{LM}}(h|e_{1},r)}{P_{w}(h|e_{1},r)}$
$=$ $\sum_{h\in\mathcal{H}}(\sum_{e_{2}\in\mathcal{E}}P(e_{2}|e_{1},h))P_{w}(h|e_{1},r)\frac{P_{\text{LM}}(h|e_{1},r)}{P_{w}(h|e_{1},r)}$
$=$ $\sum_{h\in\mathcal{H}}P_{w}(h|e_{1},r)\frac{P_{\text{LM}}(h|e_{1},r)}{P_{w}(h|e_{1},r)}$
$=$ $\text{KL}[P_{w}(\mathbf{h}|r),P_{\text{LM}}(\mathbf{h}|e_{1},r)]$
Appendix B Detailed discussion of related work
Theory on LM reasoning Many recent works are investigating LM’s reasoning ability. Geiger et al. (2021); Wu et al. (2023) aims to find the causal abstraction of an LM. (Hanna et al., 2023) tries to find circuit for year-span-prediction. Liu et al. (2023); Chi et al. (2023); Feng et al. (2023) show that CoTs enable fixed-size Transformers to perform certain types of reasoning tasks. Li et al. (2023); Razeghi et al. (2022); Wang et al. (2023) try to understand inference time in-context CoT reasoning. Our study is more relevant to the line of work analyzing the contribution of pre-training data to LM reasoning. Bi et al. (2023) analyzes how code data affect program-of-thoughts (Chen et al., 2023) reasoning ability. Xiao & Liu (2023); Zhou et al. (2023) study how reasoning length generalizes from training data. Ramesh et al. (2023) studies LMs’ compositional generalization ability. Our hypothesis also echos the conclusion of Malach (2023) that reasoning paths in training data enable supervision on intermediate steps with next-token-prediction objective, and also increase the length complexity, thus reducing time/sample complexity at training time. Prystawski et al. (2023) propose a different hypothesis that localized structure on dependencies between variables in training data is important for LM reasoning, especially CoT reasoning. Our proposed hypothesis echoes theirs and can be shown effective on more realistic data and tasks.
Logic/knowledge graph reasoning Existing methods can be divided into three categories: rule-based, GNN-based (Gori et al., 2005), and LM-based. Markov Logic Network (MLN) (Richardson & Domingos, 2006) and path ranking algorithm (PRA) (Lao et al., 2011) are two classical methods that assign weights to different logical rules. Neural Logic Programming (Yang et al., 2017) and RNN-logic (Qu et al., 2020) are two neural methods that combine the explainability of learned logical rules and the high performance of neural networks. R-GCN (Schlichtkrull et al., 2018) and NBFNet (Zhu et al., 2021) are two GNN-based methods that train a GNN on the KG and use the obtained triple embeddings. These two category methods either rely on random walks to find paths or use random walks to train GNNs. Recently, LM-based methods are shown to be highly effective on not only KG reasoning (Misra et al., 2023), but more general logical reasoning problems with text descriptions (Pan et al., 2023).
Math reasoning Recently, LM-based math reasoning models have shown to be highly effective (Azerbayev et al., 2023; Yang et al., 2023). Many works have focused on generating high-quality CoT training to improve LM’s math reasoning performance (Wang et al., 2022; Nye et al., 2022; Yuan et al., 2023). However, they all rely on the annotated Q-A pairs to generate corresponding paths with LM, which limits the size of augmented data and requires large LMs to do the CoT generation. Our proposed method does not need supervised seed data and thus can be extended to the vast amount of pre-training data. Our method is also lightweight, which only requires a small/medium LM to produce the step embeddings and then do clustering on them.
Appendix C Experiment Details
C.1 Datasets
Knowledge graph datasets For KL analysis, we focus on two KGs: Countries (Bouchard et al., 2015) and UMLS (Kok & Domingos, 2007), as they have a reasonable time complexity to compute the aggregated probabilities for long paths. The Countries (Bouchard et al., 2015) contains two relations (locatedIn and neighborOf) and 227 entities, including countries, regions, and subregions. We use the hardest version (S3) of the Countries. The Unified Medical Language System (UMLS) (Kok & Domingos, 2007) is a more complex KG built from biomedicine knowledge, containing 49 relations and 135 entities. Example entities are diseases and antibiotics, and example relations are treats and diagnoses.
We add three more datasets for computing the prediction accuracy v.s. different random walk path lengths: Kinship (Denham, 2020), NELL-995 (Xiong et al., 2017), and FB15K-237 (Toutanova et al., 2015). The Kinship dataset contains 104 entities and 26 kinship relationships among members of the Alyawarra tribe from Central Australia. The NELL-995 dataset contains 75,492 entities and 200 relations, which is built from the Web via an intelligent agent called Never-Ending Language Learner. The FB15K-237 dataset contains 14,505 entities and 237 relations derived from Freebase. We adopt a processed version of these datasets from Das et al. (2017).
Math word problem datasets We conduct experiments on three math word problem (MWP) datasets: GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021). The Grade School Math dataset (GSM8K) contains 8.5K examples of linguistically diverse grade school math world problems. The AQUA-RAT dataset contains 100K samples of mathematical problems, along with sequences of human-readable mathematical expressions in natural language. The SVAMP dataset is a testing set consisting of elementary-level MWPs. The training set is a combination of simpler MWPs: MAWPS (Koncel-Kedziorski et al., 2016) and ASDiv-A (Miao et al., 2020) with 3.5k training examples in total.
C.2 Training details
Logical reasoning We train randomly initialized GPT-2 (Radford et al., 2019) (124M parameters) with batch size 16 and learning rate 5e-4 using AdamW optimizer (Loshchilov & Hutter, 2017) on one 24G Titan GPU.
Math reasoning We continually train Llama 2 (Touvron et al., 2023b) (7B and 13B parameter versions) with batch size 16 and learning rate 2e-4 using AdamW optimizer (Loshchilov & Hutter, 2017) on one 40G A100 GPU.